# Estimating treatment effects with competing intercurrent events in randomized controlled trials

### v1

@@ 4.2 Asymptotic properties of the EIF-based estimator
We next discuss the asymptotic properties of $\hat{\tau}^{\textup{eif}}$. We first introduce additional notation. For $t\leq k$ and $a=0,1$, let $e^{*}$, $G_{a}^{*}$, $\mu_{a}^{*}$, and $S_{a}^{*}$ denote the probability limit of the estimated propensity score, estimated censoring mechanism, estimated outcome model, and estimated survival function of the AE time, respectively, i.e., $\lVert\hat{e}-e^{*}\rVert=o_{P}(1)$, $\lVert\hat{G}_{a}-G_{a}^{*}\rVert=o_{P}(1)$, $\lVert\hat{\mu}_{a}-\mu_{a}^{*}\rVert=o_{P}(1)$, and $\lVert\hat{S}_{a}-S_{a}^{*}\rVert=o_{P}(1)$. If the propensity score model is consistently estimated, we have $e^{*}=e$, and analogous results for the other three nuisance models.
In the following theorem, we provide the double robustness of the EIF estimator $\hat{\tau}^{\textup{eif}}$.
@@
Theorem 3 shows that $\hat{\tau}^{\textup{eif}}$ is consistent if at least one of the following two sets of nuisance parameters are consistently estimated: (1) the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$; (2) the propensity score model $e(X)$ and the censoring mechanism $G_{a}(t\mid X)$, for $t\leq k$ and $a=0,1$. When our data is from an RCT, i.e., the propensity score $e(X)$ is known, and thus $\hat{e}(X)$ is always correctly specified, the EIF estimator $\hat{\tau}^{\textup{eif}}$ is a consistent estimator for $\tau$ if either the censoring mechanism $G_{a}(t\mid X)$ is consistently estimated, or both the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$ are consistently estimated, for $t\leq k$ and $a=0,1$.
@@
Assume that either $\{e^{*}(X)=e(X),\allowbreak G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$ is satisfied.
@@
The convergence rate of the nuisance parameters estimation satisfy $\{||\hat{e}-e^{*}||+||\hat{G}_{a}-G_{a}^{*}||\}\{||\hat{\mu}_{a}-\mu_{a}^{*}||+||\hat{S}_{a}-S_{a}^{*}||\}=o_{P}(n^{-1/2})$ for $a=0,1$.
Assumption 3 requires either the propensity score and censoring models or the survival and outcome models to be consistently estimated. It guarantees the consistency of $\hat{\tau}^{\textup{eif}}$. Assumption 4 imposes restrictions on the nuisance model complexity and is a standard regularity condition (Van der Vaart,, 2000). Recent literature documents that we can use cross-fitting in the estimation of nuisance parameters to relax the Donsker condition (van der Laan et al.,, 2011; Chernozhukov et al.,, 2018; Díaz,, 2019). Assumption 5 imposes additional restrictions on the rate of convergence of the nuisance parameters, in addition to all of them being consistently estimated.
@@
$\hat{\tau}^{\textup{eif}}$ is a consistent and asymptotically Gaussian estimator of $\tau$ with asymptotic variance equal to $E\{D_{\tau}^{2}(O)\}$, thus achieves the semiparametric efficiency bound.
@@
Compare $\hat{\tau}^{\textup{eif}}$ with the propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$. On the one hand, $\hat{\tau}^{\textup{eif}}$ is more robust in the sense that it allows for misspecification of the censoring model and is asymptotically more efficient. Continuing the intuition discussed in Section 3.4, $\hat{\tau}^{\textup{aug}}$ only orthogonalizes the part corresponding to $A\mid X$ but not $C\mid X,A$, making it robust only to misspecification of the $A\mid X$ model, not the $C\mid X,A$ model. In contrast, $\hat{\tau}^{\textup{eif}}$ accounts for both possible sources of misspecification, leading to greater robustness and efficiency. On the other hand, $\hat{\tau}^{\textup{aug}}$ is more straightforward to implement in practice, as it only requires the estimation of $S_{a}(k\mid X)$ and $G_{a}(k\mid X)$ at time $k$ for $a=0,1$, while $\hat{\tau}^{\textup{eif}}$ requires consistent estimation of the entire survival curves $S_{a}(t\mid X)$ and $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$. We consider the choice between $\hat{\tau}^{\textup{eif}}$ and $\hat{\tau}^{\textup{aug}}$ as a trade-off between implementation simplicity and robustness and efficiency of the estimators.
5 Simulation
@@
In this section, we conduct Monte Carlo simulations to study the finite sample performance of our proposed estimators. We first generate the covariates $X=(X_{1},X_{2},X_{3})^{\tiny\textsc{t}}\in\mathbb{R}^{3}$ from three independent standard Gaussian distributions and denote $\tilde{X}_{j}=\{(X_{j}+2)^{2}-1\}/\sqrt{12}$ for $j=1,2,3$. Next, we generate the treatment assignment following $A\mid X\sim\textup{Bernoulli}(\{e(X)\})$ with $e(X)$ being the propensity score model, and generate the potential outcomes following $Y(a)\mid X\sim\mathcal{N}(\mu_{a}(X),\sigma_{a}^{2})$ with $\mu_{a}(X)$ being the outcome model for $a=0,1$. We then generate the potential values of the treatment-related ICE survival time $T(a)$ from the distribution with a survival function $S_{a}(t\mid X)=\exp[-0.002t^{1.2}\exp\{\gamma_{a}(X)\}]$ and the potential values of the treatment-unrelated ICE survival time $C(a)$ from the distributions with a survival function $G_{a}(t\mid X)=\exp[-\varrho_{a}(t)\exp\{\delta_{a}(X)\}]$, for $a=0,1$. For each of the nuisance parameters, we consider two scenarios when it will be correctly modeled and misspecified, therefore, we generate data following two different choices of each. We summarize the detailed data-generating choices in Table 2.
Table 2: Model choices for the nuisance parameters
@@
We compare the finite sample performance of the outcome estimator $\hat{\tau}^{\textup{out}}$, the weighting estimator $\hat{\tau}^{\textup{ipw}}$, the augmented weighting estimator $\hat{\tau}^{\textup{aug}}$, and the EIF estimator $\hat{\tau}^{\textup{eif}}$ based on $1000$ Monte Carlo samples with a sample size of $n=1000$. When fitting the nuisance models, we use logistic regression of $A$ on $X$ to estimate the propensity score model, the linear regression of $Y$ on $X$ on the subsample $A=a,T\wedge C>k$ to estimate the outcome model, and the Cox proportional hazard regression of $T$ and $C$ on $X$ on the subsample $A=a$ to estimate the treatment-related ICE survival model and the censoring model, respectively, for $a=0,1$. We use $X$ in all regression models, therefore, when the true data-generating processes involve non-linear functions of $X$, the fitted nuisance models suffer from misspecification. We evaluate the performance of the four estimators with reported bias, standard deviation, and coverage probability using a nonparametric bootstrap with $200$ bootstrap samples in Table 3.
Table 3: Finite sample performance of the four estimators. For each estimator, we report the finite-sample bias, standard deviation (SD), and the coverage rate (CR) of a 95% confidence interval, which is constructed using a nonparametric bootstrap with 200 bootstrap iterations. Each row corresponds to a different data-generating regime.

### v2

@@ 4.2 Asymptotic properties of the EIF-based estimator
We next discuss the asymptotic properties of $\hat{\tau}^{\textup{eif}}$. We first introduce additional notation. For $t\leq k$ and $a=0,1$, let $e^{*}$, $G_{a}^{*}$, $\mu_{a}^{*}$, and $S_{a}^{*}$ denote the probability limit of the estimated nuisance functions $\hat{e}$, $\hat{G}_{a}$, $\hat{\mu}_{a}$, $\hat{S}_{a}$, respectively, i.e., $\lVert\hat{e}-e^{*}\rVert=o_{P}(1)$, $\lVert\hat{G}_{a}-G_{a}^{*}\rVert=o_{P}(1)$, $\lVert\hat{\mu}_{a}-\mu_{a}^{*}\rVert=o_{P}(1)$, and $\lVert\hat{S}_{a}-S_{a}^{*}\rVert=o_{P}(1)$. If a given nuisance model is correctly specified, its corresponding limit equals the true function. For example, if the propensity score model is consistent, then $e^{*}=e$, and similar results hold for the other three nuisance components.
In the following theorem, we provide the double robustness of the EIF-based estimator $\hat{\tau}^{\textup{eif}}$.
@@
Theorem 3 shows that $\hat{\tau}^{\textup{eif}}$ is consistent if at least one of the following two sets of nuisance parameters are consistently estimated: (1) the outcome model $\mu_{a}(X)$ and the survival function for the treatment-related ICE $S_{a}(t\mid X)$; (2) the propensity score model $e(X)$ and the survival function for the treatment-unrelated ICE $G_{a}(t\mid X)$, for $t\leq k$ and $a=0,1$. When our data is from an RCT, i.e., the propensity score $e(X)$ is known by design, so $\hat{e}(X)$ can be correctly specified. Consequently, the EIF-based estimator $\hat{\tau}^{\textup{eif}}$ is a consistent estimator for $\tau$ if either $G_{a}(t\mid X)$ is consistently estimated, or both $\mu_{a}(X)$ and $S_{a}(t\mid X)$ are consistently estimated, for $t\leq k$ and $a=0,1$.
@@
Assume that either $\{e^{*}(X)=e(X),G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$ is satisfied.
@@
The convergence rate of the nuisance parameters estimation satisfies $\{||\hat{e}-e^{*}||+||\hat{G}_{a}-G_{a}^{*}||\}\{||\hat{\mu}_{a}-\mu_{a}^{*}||+||\hat{S}_{a}-S_{a}^{*}||\}=o_{P}(n^{-1/2})$ for $a=0,1$.
Assumption 3 requires either the propensity score and survival function of the treatment-unrelated ICE or the outcome and survival function of the treatment-related ICE to be consistently estimated. It guarantees the consistency of $\hat{\tau}^{\textup{eif}}$. Assumption 4 imposes restrictions on the nuisance model complexity and is a standard regularity condition (Van der Vaart,, 2000). We can also employ flexible machine learning models with cross-fitting techniques in the estimation of nuisance parameters to relax the Donsker condition (Pfanzagl and Wefelmeyer,, 1985; Klaassen,, 1987; Zheng and van der Laan,, 2011; Chernozhukov et al.,, 2018). Finally, Assumption 5 imposes additional restrictions on the rate of convergence of the nuisance parameters, in addition to all of them being consistently estimated.
@@
The estimator $\hat{\tau}^{\textup{eif}}$ is a consistent and asymptotically Normal estimator of $\tau$ with asymptotic variance equal to $E\{D_{\tau}^{2}(O)\}$, thus achieving the semiparametric efficiency bound.
@@
4.3 Comparing the estimators
In real data analysis, we recommend implementing and comparing all four proposed estimators to assess whether they provide coherent scientific implications. Theoretical comparisons among these estimators primarily focus on two key dimensions: robustness to nuisance model misspecification and asymptotic efficiency. In practice, examining how the point estimates differ across these estimators can provide insight into the presence of nuisance model misspecification. Specifically, certain pairwise differences among the estimators can serve as informal diagnostics. For example, if $e(X)$ is correctly specified, the difference between $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ should converge to zero. Similarly, if $\mu_{a}(X)$ and $S_{a}(k\mid X)$ are correctly specified, the difference between $\hat{\tau}^{\textup{out}}$ and $\hat{\tau}^{\textup{aug}}$ should converge to zero. Finally, if $G_{a}(t\mid X)$ is correctly specified, the difference between $\hat{\tau}^{\textup{eif}}$ and $\hat{\tau}^{\textup{aug}}$ should converge to zero. Therefore, when implementing all four estimators in practice, the observed difference among them may serve as an informal diagnostic tool to detect possible nuisance model misspecification. We provide more details in Section A.1 in the supplementary material.
@@ 4.3 Comparing the estimators
Among the four estimators, a key comparison lies between the EIF-based estimator $\hat{\tau}^{\textup{eif}}$ with the augmented weighting estimator $\hat{\tau}^{\textup{aug}}$. On the one hand, $\hat{\tau}^{\textup{eif}}$ is more robust and asymptotically efficient: it allows for misspecification of the censoring model $C\mid A,X$ and it achieves the semiparametric efficiency bound under correct nuisance specification. Continuing the intuition discussed in Section 3.4, $\hat{\tau}^{\textup{aug}}$ only orthogonalizes the part corresponding to $A\mid X$ but not $C\mid X,A$, making it robust only to misspecification of the $A\mid X$ model, not the $C\mid X,A$ model. In contrast, $\hat{\tau}^{\textup{eif}}$ accounts for both possible sources of misspecification, leading to greater robustness and efficiency. On the other hand, $\hat{\tau}^{\textup{aug}}$ is more straightforward to implement in practice, as it only requires the estimation of $S_{a}(k\mid X)$ and $G_{a}(k\mid X)$ at time $k$ for $a=0,1$, while $\hat{\tau}^{\textup{eif}}$ requires consistent estimation of the entire survival curves $S_{a}(t\mid X)$ and $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$. We consider the choice between $\hat{\tau}^{\textup{eif}}$ and $\hat{\tau}^{\textup{aug}}$ as a trade-off between implementation simplicity and robustness and efficiency of the estimators.
Finally, we note the theoretical possibility of applying a martingale-based correction to the basic weighting estimator $\hat{\tau}^{\textup{ipw}}$, in a manner analogous to the correction in the EIF-based estimator $\hat{\tau}^{\textup{eif}}$ compared with $\hat{\tau}^{\textup{aug}}$. Such a corrected estimator would be robust to misspecification of $G_{a}(t\mid X)$. However, it remains less attractive in practice: it is neither robust to misspecification of the propensity score nor asymptotically efficient, and it lacks the implementation simplicity of $\hat{\tau}^{\textup{aug}}$. For these reasons, we do not recommend it, though we include its explicit form in Section A.2 of the supplementary material only for theoretical completeness.
4.4 Simulation results
We conduct Monte Carlo simulations to evaluate the finite-sample performance of four estimators under various combinations of correctly specified and misspecified nuisance models. Consistent with our theory, $\hat{\tau}^{\textup{eif}}$ demonstrates the greatest robustness across all misspecification scenarios, maintaining low bias and valid coverage even when multiple nuisance models are incorrect. We relegate the detailed results to Section B in the supplementary material.
5 Analyzing two phase-3 trials
5.1 Data analysis
In this section, we re-analyze data from two trials on systemic lupus erythematosus (Morand et al.,, 2023; Petri et al.,, 2023). Both trials are double-blinded, randomized, placebo-controlled phase-3 trials. There are three treatment arms: 2mg baritinib, 4mg baritinib, and placebo. The primary outcome in both studies is the Systemic lupus erythematosus Responder Index 4 (SRI4) at week 52, a binary composite endpoint that reflects clinical response. A participant is classified as a responder if they show a predefined improvement in disease activity, without overall worsening or the emergence of significant disease activity in new organ systems. Baseline covariates used for model fitting include geographic region, baseline corticosteroid use, and the Physician’s Global Assessment score measured at baseline.
A substantial proportion of participants in both trials experienced ICEs: 218 out of 760 in one trial and 211 out of 775 in the other. As described in Section 1.2, we classify these ICEs into two categories: treatment-related ICEs (82.6% and 84.4% in the two trials, respectively) and treatment-unrelated ICEs (17.4% and 15.6%). We conduct separate comparisons of the 2mg and 4mg baricitinib treatment arms versus the placebo group. For each comparison, we apply all four estimators: $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, $\hat{\tau}^{\textup{aug}}$, and $\hat{\tau}^{\textup{eif}}$. We use logistic regression to estimate the propensity score and outcome models, and use Cox proportional hazards regression to separately estimate the survival functions for the two types of ICEs. We conduct variance estimation and inference using a nonparametric bootstrap with 500 replicates. Table 2 presents the resulting point estimates, their estimated standard errors, and corresponding $p$-values.
Table 2: Treatment effect of Baricitinib on the primary outcome SRI4. The first four columns report the results of Trial 1 and the last four columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the composite outcome average treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on all four estimators.
Trial 1 (Petri et al.,, 2023) Trial 2 (Morand et al.,, 2023)
2mg point 0.030 0.029 0.026 0.026 0.019 0.019 0.022 0.022
se 0.042 0.043 0.042 0.043 0.042 0.043 0.043 0.042
$p$-value 0.479 0.504 0.534 0.540 0.643 0.662 0.602 0.606
4mg point 0.113 0.120 0.115 0.113 $-0.002$ $-0.002$ $-0.002$ $-0.002$
se 0.046 0.046 0.046 0.046 0.042 0.042 0.042 0.042
$p$-value 0.013 0.008 0.012 0.013 0.961 0.962 0.966 0.969
5.2 Interpretation of the statistical results
Across both trials and treatment arms, the point estimates and $p$-values from the four methods are generally consistent. Our results suggest a statistically significant positive effect of 4mg baricitinib on the SRI4 outcome in Trial 1, while in Trial 2, the effect is not significant and the point estimate is slightly negative. No significant treatment effect is observed for the 2 mg baricitinib dose in either trial. As discussed in Section 4.3, though not a formal statistical test, qualitatively, the robustness of estimates across the four estimators in Table 2 provides empirical evidence that our nuisance models may not be severely misspecified.
5.3 Comparison with ad-hoc methods
Next, we use the two commonly used ad-hoc methods and compare them with our proposed methods. The first method is the non-responder imputation method that assigns an outcome as 00 whenever an ICE occurs. It is equivalent to using the composite outcome strategy that treats all ICEs as treatment-related, leading to a composite outcome equal to 00. The second method naively applies a hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on observed covariates.
Table 3 reports the estimated treatment effect of both treatment arms across the two trials using the inverse probability weighting estimator. The non-responder imputation method generally yields slightly smaller effect estimates than our proposed methods, although the differences are modest, likely due to the relatively low proportion of treatment-unrelated ICEs in both trials. Nevertheless, our proposed approach targets a clinically more meaningful causal estimand. The second ad hoc method naively applies the hypothetical strategy to all ICEs, assuming that their occurrence is conditionally independent of the potential outcomes given baseline covariates. This assumption is untestable and likely violated in practice. For example, treatment discontinuation due to adverse events or lack of efficacy is plausibly related to a patient’s potential outcome. As shown in Table 3, this method produces treatment effect estimates that differ substantially from those obtained using our proposed approach, potentially leading to misleading clinical conclusions. These empirical findings are consistent with the large-sample bias observed in the simulation study and underscore the importance of using appropriate strategies tailored to different types of ICEs.
Table 3: Treatment effect of Baricitinib on the primary outcome SRI4 using non-responder imputation (NRI) and hypothetical strategy (HS). The first two columns report the results of Trial 1 and the last two columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the estimated treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on the inverse probability-weighting estimator.
Trial 1 Trial 2
NRI HS NRI HS
2mg point 0.039 0.011 0.010 0.022
se 0.043 0.049 0.042 0.050
$p$-value 0.362 0.815 0.817 0.654
4mg point 0.093 0.073 $-0.008$ 0.044
se 0.044 0.050 0.041 0.051
$p$-value 0.036 0.147 0.847 0.382
6 Future directions
In this study, we address the challenges posed by ICEs in RCTs by proposing methods to handle competing ICEs. We classify ICEs into treatment-related and treatment-unrelated events and apply different strategies to identify a clinically meaningful causal effect. For treatment-related ICEs, which are often informative about a patient’s outcome, we use a composite variable strategy that assigns an outcome value indicative of treatment failure. For treatment-unrelated ICEs, we apply a hypothetical strategy, assuming their timing is conditionally independent of the outcome given treatment and baseline covariates, and envisioning a scenario in which such events do not occur. The central thesis of this paper is to address the challenge of competing ICEs, where the first ICE censors all subsequent ones. In this paper, we propose a principled framework that carefully formulates the estimand, establishes its nonparametric identification and semiparametric estimation theory, and introduces weighting, outcome regression, and doubly robust estimators. While our proposed framework provides a rigorous and flexible approach for handling competing ICEs in RCTs, several challenges and extensions remain open for future research.
6.1 Data collection and ICE classification
Our proposed methods have broad applicability across various therapeutic areas, including immunology, oncology, and cardiology, where treatment discontinuation and other ICEs frequently occur. However, the effectiveness of these approaches relies on the accurate classification of ICEs, which should be performed in collaboration with clinicians to ensure clinical relevance in trial analyses. To support this, a modernized case report form is needed to enable more granular documentation of the timing, reasons, and magnitude of treatment discontinuation. For example, the reason for the event should be specified, such as discontinuation due to toxicity versus lack of efficacy. In some cases, the event may need to meet a threshold of magnitude, such as the use of additional medication exceeding a specified duration or dose. Additionally, the timing of the event may be relevant, particularly in relation to its proximity to outcome assessment. Fortunately, a cross-industry PHUSE working group is tackling this problem, and a recommended new case report form will be available soon (PHUSE Working Group,, 2024).
6.2 Random $K$
In practice, the measurement of the outcome of interest $Y$ does not necessarily happen at a fixed point $k$. For instance, the time $K$ that a patient visits the clinic and takes the measurement can be treated as a random variable independent of treatment $A$, potential outcomes $Y(a)$, and both types of ICEs $T(a)$ and $C(a)$ for $a=0,1$. For a random $K$, the analogous formulas to both the previous identification formulas (2) and (3) hold, with a replacement of $k$ to $K$, i.e.,
If $K$ is a pre-treatment covariate that is observed, the proposed four estimators carry over by replacing the fixed $k$ with the observed values of $K$. If $K$ is also treated as a post-treatment variable, we can construct a weighting estimator following (10). However, (9) no longer provides a feasible identification formula unless $S_{a}(K\mid X)$ is identified for $a=0,1$. Considering the presence of three competing time-to-event random variables, it becomes necessary to employ competing risks models. We defer this analysis to future work.
Acknowledgement
Sizhu Lu and Peng Ding were partially supported by the U.S. National Science Foundation grants # 1745640 and #1945136. Ting Ye was partially supported by the HIV Prevention Trials Network (HPTN) and NIH grant: NIAID 5 UM1 AI068617.
Supplementary material
The supplementary material includes additional technical details, simulation results, and proofs of all theorems and propositions.
Supplementary Material
Section A provides additional results on comparing the four proposed estimators and another martingale-corrected weighting estimator, discussed in Section 4.3 of the main paper.
Section B reports detailed results from Monte Carlo simulations that assess the finite-sample performance of our proposed estimators.
Section C provides proofs of all theorems and propositions.
Appendix A Additional results
A.1 Pairwise differences between the estimators
[…]
These probability limits indicate the impact of nuisance model misspecification on the pairwise comparisons. For example, if $G_{1}(t\mid X)$ is correctly specified, the expected difference between $\hat{\mu}_{1}^{\textup{eif}}$ and $\hat{\mu}_{1}^{\textup{aug}}$ should be small. Similar logic applies to other comparisons. Therefore, when implementing all four estimators in practice, the observed difference among them may serve as an informal diagnostic tool to detect possible nuisance model misspecification.
A.2 The martingale-corrected weighting estimator
In this subsection, we present the explicit form of the martingale-corrected weighting estimator introduced in Section 4.3. Drawing from the proof of Theorem 2 in Section C.3, if we do not project onto the propensity score nuisance tangent space in Step 2, the following identification formula for $\mu_{1}$ holds:
$\mu_{1}$ $=$ $E\left[\frac{A}{e(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}(k\mid X)}+\mu_{1}(X)S_{1}(k\mid X)\int_{0}^{\tilde{T}}\frac{\mathrm{~{}d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}\right],$
which motivates the following estimator for $\mu_{1}$:
$\hat{\mu}_{1}^{\textup{mc-ipw}}$ $=$ $n^{-1}\sum_{i=1}^{n}\frac{A_{i}Y_{i}1(T_{i}\wedge C_{i}>k)}{\hat{e}(X_{i})\hat{G}_{1}(k\mid X_{i})}+n^{-1}\sum_{i=1}^{n}\frac{A_{i}}{\hat{e}(X_{i})}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}.$
The control counterpart, $\hat{\mu}_{0}^{\textup{mc-ipw}}$, is defined analogously, and the estimator for $\tau$ is given by $\hat{\tau}^{\textup{mc-ipw}}=\hat{\mu}_{1}^{\textup{mc-ipw}}-\hat{\mu}_{0}^{\textup{mc-ipw}}$.
Under Assumuptions 1 and 2, and assuming correct specification of $e(X)$, the estimator $\hat{\tau}^{\textup{mc-ipw}}$ is consistent for $\tau$ if either $G_{a}(t\mid X)$ is correct, or both $\mu_{a}(X)$ and $S_{a}(t\mid X)$ are correct for $t\leq k$ and $a=0,1$. Although the martingale correction enhances robustness to misspecification of $G_{a}(t\mid X)$, $\hat{\tau}^{\textup{mc-ipw}}$ is still less attractive than $\hat{\tau}^{\textup{eif}}$: it does not achieve the efficiency bound and is not robust to misspecification of $e(X)$. Moreover, it does not inherit the implementation simplicity of $\hat{\tau}^{\textup{aug}}$. In particular, it requires estimating the full survival curves for both the treatment-unrelated ICEs $G_{a}(t\mid X)$ and treatment-related ICEs $S_{a}(t\mid X)$ over $t\leq k$, making it computationally demanding without offering clear practical advantages.
Appendix B Simulation
In this section, we conduct Monte Carlo simulations to study the finite sample performance of our proposed estimators.
B.1 Data generating processes
@@
We first generate the covariates $X=(X_{1},X_{2},X_{3})^{\tiny\textsc{t}}\in\mathbb{R}^{3}$ from three independent standard Gaussian distributions and denote $\tilde{X}_{j}=\{(X_{j}+2)^{2}-1\}/\sqrt{12}$ for $j=1,2,3$. Next, we generate the treatment assignment following $A\mid X\sim\textup{Bernoulli}(\{e(X)\})$ with $e(X)$ being the propensity score model, and generate the potential outcomes following $Y(a)\mid X\sim\mathcal{N}(\mu_{a}(X),\sigma_{a}^{2})$ with $\mu_{a}(X)$ being the outcome model for $a=0,1$. We then generate the potential values of the treatment-related ICE survival time $T(a)$ from the distribution with a survival function
$S_{a}(t\mid X)=\exp[-0.002t^{1.2}\exp\{\gamma_{a}(X)\}]$
and the potential values of the treatment-unrelated ICE survival time $C(a)$ from the distributions with a survival function
$G_{a}(t\mid X)=\exp[-\varrho_{a}(t)\exp\{\delta_{a}(X)\}],$
for $a=0,1$. For each of the nuisance parameters, we consider two scenarios when it will be correctly modeled and misspecified, therefore, we generate data following two different choices of each. We summarize the detailed data-generating choices in Table S1.
Table S1: Model choices for the nuisance parameters
@@
B.2 Simulation results
We compare the finite sample performance of the outcome estimator $\hat{\tau}^{\textup{out}}$, the weighting estimator $\hat{\tau}^{\textup{ipw}}$, the augmented weighting estimator $\hat{\tau}^{\textup{aug}}$, and the EIF estimator $\hat{\tau}^{\textup{eif}}$ based on $1000$ Monte Carlo samples with a sample size of $n=1000$. When fitting the nuisance models, we use logistic regression of $A$ on $X$ to estimate the propensity score model, the linear regression of $Y$ on $X$ on the subsample $A=a,T\wedge C>k$ to estimate the outcome model, and the Cox proportional hazard regression of $T$ and $C$ on $X$ on the subsample $A=a$ to estimate the treatment-related ICE survival model and the censoring model, respectively, for $a=0,1$. We use $X$ in all regression models, therefore, when the true data-generating processes involve non-linear functions of $X$, the fitted nuisance models suffer from misspecification. We evaluate the performance of the four estimators with reported bias, standard deviation, and coverage probability using a nonparametric bootstrap with $200$ bootstrap samples in Table S2.
Table S2: Finite sample performance of the four estimators. For each estimator, we report the finite-sample bias, standard deviation (SD), and the coverage rate (CR) of a 95% confidence interval, which is constructed using a nonparametric bootstrap with 200 bootstrap iterations. Each row corresponds to a different data-generating regime.